What is the measure of an angle if the measure of its compliment is 10 degrees more than twice its measure? (Show your solution)
Answer by mananth(16946)

let the measure of angle be x
thecomplement will be 90-x
...
90-x = 2x+10
90-10=2x+x
80 = 3x
x= 80/3
x= 26.6 deg
the complement = 90-x= 63.4 deg
...
m.ananth@hotmail.ca